Solve the system by graphing.
3x + y = 6
3x – y = 0

Answer by Lacey020991(49)

a) 3x + y = 6
b) 3x – y = 0
.
1) put each equation into y = mx + b form to graph
.
3x + y = 6
y = -3x + 6
.
3x - y = 0
-y = -3x
y = 3x
.
.
.
Photo Sharing and Video Hosting at Photobucket
.
The solution would be (1,3)
.
.
Hope this helped you!